# Hefei, China OOD Detection for AV-GZSL: A General Framework

### v1

@@ 1 Introduction
[…]
@@ 3.1 Feature Generator
Embedding loss. We organize the real and synthesized features as matched (same classes) and unmatched (different classes) pairs. Then, a Cosine Embedding loss $\mathbf{CE}$ is employed to minimize and maximize the distance between the matched and unmatched features, respectively. Thus, $\mathcal{L}_{emb}=\mathbf{CE}(x,\tilde{x})$.
@@ 3.2 Out-Of-Distribution Detector
$\mathcal{L}_{ent}=\mathbf{H}\left(O_{ent}(x)\right)-\mathbf{H}\left(O_{ent}(\tilde{x})\right),$ (4)
@@ 3.4 Unseen Classifier
[…]
@@
Thus, the total loss of the unseen classifier is formulated as $\mathcal{L}_{uc}=\mathcal{L}^{+}+\mathcal{L}^{-}$, where ±$\mathcal{L}^{\pm}=\mathcal{L}_{trip}^{\pm}+\mathcal{L}_{rec}^{\pm}+\mathcal{L}_{reg}^{\pm}$.

### v2

@@ 1 Introduction
Recently, several studies [Parida et al.(2020)Parida, Matiyali, Guha, and Sharma, Mazumder et al.(2021)Mazumder, Singh, Parida, and Namboodiri] have delved into the field of AV-GZSL, innovating methods that project both modalities into a common embedding space and measure distances from class label text embeddings. $\bmvaOneDot[Mercea et al$.(2022b)Mercea, Riesch, Koepke, and Akata] introduced a integrated framework, employing a transformer-based cross-attention mechanism on temporally averaged audio and visual input features. Subsequently, numerous subsequent works have emerged, such as [Mercea et al.(2022a)Mercea, Hummel, Koepke, and Akata, Li et al.(2023c)Li, Luo, and Du, Li et al.(2023a)Li, Ma, Deng, Man, and Fan, Hong et al.(2023)Hong, Hayder, Han, Fang, Harandi, , and Petersson, Li et al.(2023b)Li, Zhao, Ma, Wang, Fan, , and Tian], including a partially generative approach [Zheng et al.(2023)Zheng, Hong, and Farazi] which synthesizes unseen features as negative samples in contrastive loss. However, all these approaches are essentially embedding-based and suffer significantly from the domain shift problem [Pourpanah et al.(2022)Pourpanah, Abdar, Luo, Zhou, Wang, Lim, and Wu]. Although they employed calibrated stacking [Chao et al.(2016)Chao, Changpinyo, Gong, and Sha] to reduce the bias, it is merely a temporary solution with limited effectiveness.
@@ 3.1 Feature Generator
[…]
@@ 3.2 Out-Of-Distribution Detector
[…]
@@ 3.4 Unseen Classifier
where $\mathbf{T}(\cdot)$ represents the triplet loss function $\mathbf{T}(x,y,z)=\max\{\left\lVert x-y\right\rVert_{2}-\left\lVert x-z\right\rVert_{2}+\mu,0\}$. Here, $x$, $y$, and $z$ are anchor embeddings, and $\mu$ denotes the margin. $\mathcal{O}_{m}^{proj,\pm}$ represents the output from the projector in the joint space, where $m\in\{a,v,t\}$ denotes three modalities. Superscripts $\{+,-\}$ represent positive and negative samples, respectively. These four types of triplet loss functions cooperate to align contrastive audio, visual, and textual label embeddings.
@@
Thus, the total loss of the unseen classifier is formulated as $\mathcal{L}_{uc}=\mathcal{L}^{+}+\mathcal{L}^{-}$, where $\mathcal{L}^{\pm}=\mathcal{L}_{trip}^{\pm}+\mathcal{L}_{rec}^{\pm}+\mathcal{L}_{reg}^{\pm}$.